Subject: APO 3.5m users committee minutes, May 2012

From: Suzanne Hawley

Submitted: Wed, 30 May 2012 12:28:49 -0700 (PDT)

Message number: 1273 (previous: 1272, next: 1274 up: Index)

 		APO 3.5-m Users Committee Phonecon, 5/07/12

Attending:  Suzanne Hawley, Mark Klaene, Scott Anderson, Bill Ketzeback, 
Sean Moran, Jon Holtzman, John Bally, Ed Turner (for Michael Strauss), and 
Bruce Gillespie

Absent:  Remy Indebetouw, Al Harper


User feedback and comments from institutional representatives:

  o Chicago (Al Harper) - No report.

  o Washington (Scott Anderson) - Scott had nothing to report.

  o New Mexico State (Jon Holtzman) - Jon had nothing to report.

  o Princeton (Ed Turner) - Ed had nothing to report.

  o Colorado (John Bally) - John had nothing to report.

  o Virginia (Remy Indebetouw) - No report.

  o Johns Hopkins (Sean Moran) - Sean had nothing to report.


Discussion of telescope/instruments report:

Mark announced that we are again experiencing high fire-danger conditions 
at APO, and the annual Miller moth invasion has begun and appears to be 
heralding a "bad" moth year.  The telescope has been trouble-free.  On 
instruments, the DIS blue camera developed scattered light problems. We 
will perform a major service on DIS this summer, and have pumped on the 
dewar in the meantime, resulting in some improvement.  We fixed a 
cable-snagging problem with Agile.  SPIcam's vacuum problem got recently 
worse, so it was opened for major service and will be returned to 
operations by the end of May.  The masks for TripleSpec were repaired, but 
this needs finishing work in TUI.  The new grating will get installed in 
the echelle during the summer shutdown. We are working to insure that 
there are baseline throughput data for comparison with the new grating? 
The situation with NIC-FPS was discussed below as a separate topic.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                 04/05/12 through 05/02/12

0) Overview

Telescope continued to run well, and the overall good weather continued. 
The fire danger finally hit "High" however the overall mood in the heavy 
forested regions is that the risk of a large forest fire is still small. 
The miller moth season does appear to be forming.  Large infestations of 
moths often cause telescope drive slippage and increased spotting on 

1) Telescope

The telescope has been operational with no issues of note.

2) Instruments

DIS Blue continues to suffer from vacuum problems including some increased 
scattered light resulting from contamination.  We are working on the 
vacuum and the blue side of DIS is out of service this week while we pump 
on it.  Significant dewar work is planned for the summer and we are 
focusing on keeping it operational until then.

Agile had an issue with the rotator.  The problem was traced to a 
combination of rotation angle and altitude that caused the umbilical to 
hang up.  Problem was corrected the next day.

SPIcam vacuum continued to worsen and the instrument was taken off line at 
the completion of scheduled observing at the end of April.  Vacuum and 
coldhead work has progressed smoothly.  The coldhead has been replaced, 
o-rings have been serviced, several vacuum connections have been 
eliminated, a new vacuum gauge installed and a loose nylon screw removed 
from inside the dewar.  We have started looking at the shutter and found 
substantial wear on the shutter blade caused by interference with the 
housing. Additionally the motor-gearbox assembly was leaking oil.  After 
the shutter and final vacuum pump down is completed, the cryotiger 
compressor will be serviced and the instrument reassembled.  We plan to 
have the instrument operational in time for scheduled observing to resume 
at the end of May.

TSpec maskfile has been remade and tested at various binning factors. 
Data are being analyzed.

Echelle is operational.  Plan is for installation of the new grating 
during summer shutdown.  The new ecam has been running much smoother.

NIC-FPS has been operating in shared-risk observing status.  The image 
corruption issues are still present on occasion but have not been reported 
to be affecting science observations.  We have received a proposal from 
UVa to address NICFPS issues.  There is a separate agenda/action item to 
discuss this proposal.


UVa NIC-FPS proposal, NIC-FPS action item:

Suzanne said that the Users Committee members had an action item to ask 
their users about the impact of taking NIC-FPS out of service for six 
months so that it can be repaired at UVa.  All five of the committee 
members present said that they asked, and heard back that there would be 
no significant impact.  Scott added that there was concern about potential 
risks to the instrument.  The instrument team at UVa (Mike Skrutskie, John 
Wilson, Fred Hearty, Matt Nelson) submitted a proposal to repair the 
electronic problem(s) with NIC-FPS, and assess what it would take to 
upgrade the camera to a 2048x2048 detector and get the NIR etalon to work. 
Their rough cost estimate for this work is $100k.  Given that NIC-FPS is 
currently usable, the committee felt that it would be of marginal value to 
users to make these repairs and/or upgrades.  Suzanne said that she would 
contact UChicago users directly, and that any other users who have 
opinions on this matter should contact her directly.  This closes the 
pending actions item.


Dewpoint sensors:

Bill reminded us that a new dewpoint sensor was purchased and installed 
over a year ago.  It generally works more accurately than the old dewpoint 
sensor we've used for many years, but requires a workaround for a built-in 
calibration feature that occurs during the night.  We are working with the 
manufacturer, and are also planning to buy a second sensor for use inside 
the enclosure.  If all goes as hoped, the new sensors should enable us to 
relax the dewpoint closure limit somewhat, which could allow us a few more 
nights on the sky per year.


Summer shutdown timing and work:

Mark reported that the shutdown will run from 23 July to August 9, with 
the plan being to be back on the sky the night of 10 August, weather 
permitting.  The primary mirror will be washed, and M2 and M3 will be 
realuminized.  The rest of the shutdown will involve work on the 
instruments, including DIS, echelle, axis controllers, cryotiger 
maintenance, etc.


ARCSAT upgrade:

Suzanne announced that Dmitry Bizyaev received a grant to buy a new camera 
for ARCSAT which is presently being installed. The UW high-speed camera 
(Flarecam) is also being used at the telescope.  The plan is that these 
two cameras will be generally available for use on ARCST within the next 
few months.  We are working on a control system for remote operations of 
the telescope, but at present it is not robust enough for general use. 
When it is ready, we hope that ARCSAT will be an attractive facility for 
present, and possibly new, ARC partners.  The funding plan and proposal 
process for ARCSAT time remains to be developed once it is operational. 
Suzanne is interested in hearing from the ARC user community about their 
interest in using ARCSAT.


Advertising for new ARC (APO 3.5m) partners:

Suzanne reminded the committee that a couple of institutions are looking 
to sell their ARC shares.  It is also the case that other mid-sized 
telescopes are looking for new partners, so we need to be acting 
expeditiously.  There is an ARC Partners Committee that has been working 
on a strategy to search for new ARC partners.  A URL has been set up at 
APO that contains information on the 3.5-m telescope, its instrumentation, 
and scientific highlights.  We are sending letters to select institutions, 
offering an invitation to discuss possible relationships with ARC for 
access to the 3.5-m telescope, either as full ARC partners, on a leasing 
basis, or as part of an SDSS-like Participation Group.  Bruce will forward 
the URL to the Users Committee members.  If anyone knows of an 
institution/contact that should be approached, please notify Suzanne, Rene 
Walterbos, and Bruce.



[open from previous meeting]:

===> ACTION: Users committee members to discuss implications of taking 
NICFPS out of service for six months.  In particular, are users happy with 
the current state, is it worth trying to fix?

 	Status:  Closed

===> ACTION: Send new partner suggestions to Suzanne, Rene and Bruce.

 	Status:  Open

[new actions from this meeting]:



Next meeting:  The next Users Committee phonecon will be on Monday 4 June 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to 
the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1273 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to